OpenStudy (anonymous):

The length of a rectangle is 4 feet more than twice the width. The perimeter is 116 feet. Find the dimensions of the rectangle. Please explain

OpenStudy (anonymous):

Sorry what was dimensions again?

OpenStudy (anonymous):

 since perimeter is P=2L + 2W = 2(2W+4) + 2W, solve the equation to find W: 116 = 2(2W+4) + 2W